# HyperSBINN: A Hypernetwork-Enhanced Systems Biology-Informed Neural Network for Efficient Drug Cardiosafety Assessment

### v1

@@ 6.1 Performance of the hyperSBINN model
This section presents a series of experimental results. Figure 2 presents the convergence behavior of the loss components described in Section 3.1 during the training of the HyperSBINN model. The graph demonstrates the training dynamics of the HyperSBINN, with a significant initial drop in the dynamic loss ($\mathcal{L}^{\text{ODE}}$) indicating rapid learning of the underlying ODEs. The steady decrease and eventual stabilization of the initial condition loss and observation loss suggest effective learning of initial conditions and a good fit to the observed data. One can notice high variability of the ODE loss, which makes the learning highly sensitive to the learning rate schedule. The training of the HyperSBINN takes around two hours but once it is trained, an evaluation takes only a few seconds which is twenty times faster than using the ODE solver.
@@
The results from the HyperSBINN model presented in 3 closely match those obtained from the numerical solver across all variables, with minor differences observed in transition phases. This close alignment indicates that the HyperSBINN effectively captures the system’s dynamics and provides comparable performance to the traditional solver. The slight discrepancies in signal with low range of variations (such as $h_{\text{K-S}}$, $h_{\text{Ca-L}}$ or $h_{\text{To-S}}$ might be corrected by adjustment of the different ODE weights.
Figure 3 also presents a comparative analysis between the traditional solver and the HyperSBINN model for two sets of parameterizations over a time span of 500 milliseconds. The HyperSBINN model demonstrates high accuracy in replicating the results from the traditional solver across all variables. This can be seen in Table 1 which provides the normalized squared differences between curves outputted by the numerical solver and curves provided by the HyperSBINN on a test set of drug configurations. The order of magnitude is low for a vast majority of configurations. However, a few configurations (two over 500 in our test set) present a strong discrepancy.
We have noticed the importance of observed data, as the training without the observation loss would lead, in most of the time to solution which are constant outside of the initial conditions (not shown here). Therefore, to ensure the model captures the full range of possible solutions, it is crucial to incorporate additional data points that represent the solutions of the ODE system at different time points $t$, beyond just the initial conditions and ODE loss.
Curve Min Q1 Median Mean Q3 Max
$V(t)$ 0.01 0.07 0.16 1.71 0.59 161.49
$f_{\text{Ca-L}}$ 0.02 0.05 0.10 1.21 0.42 173.76
$h_{\text{Ca-L}}$ 0.08 0.23 0.54 2.36 1.41 178.18
$f_{\text{Na-F}}$ 0.19 0.27 0.46 4.61 1.61 436.78
$f_{\text{Na-L}}$ 0.12 0.30 0.42 0.99 0.72 43.92
$h_{\text{Na-f}}$ 0.00 0.03 0.10 1.46 0.55 163.84
$h_{\text{Na-L}}$ 0.01 0.03 0.08 1.26 0.40 185.51
$f_{\text{To-F}}$ 0.03 0.05 0.10 1.04 0.30 228.93
$f_{\text{To-S}}$ 0.07 0.11 0.15 1.25 0.35 250.48
$h_{\text{To-F}}$ 0.11 0.16 0.20 1.24 0.41 185.34
$h_{\text{To-S}}$ 7.40 13.36 14.67 17.33 16.83 219.86
$f_{\text{K-R}}$ 0.04 0.09 0.15 1.40 0.40 183.74
$f_{\text{K-S}}$ 0.10 0.23 0.47 1.83 1.17 193.82
$h_{\text{K-S}}$ 0.92 1.55 1.84 3.37 2.97 144.10
Table 1: Normalized squared differences between curves outputted by the solver and curves outputted by the SBINN. All metrics were multiplied by 1000, for instance, a value of 200 corresponds to a normalized squared difference of 0.2.
Figure 3: Comparison of the solver and hyperPINN curves for 2 different drug configurations never seen on the training set (one color per configuration) for each 14 solutions of the biological system. Significance of all signals names is depicted in Appendix. Our main signal of interest is $V(t)$ (top left).
@@ 6.2 Learning specific biomarker of the CAP
Table 2 shows the absolute error of prediction of APD90 when this quantity is computed using the voltage curve output by the hyperSBINN. Overall, the prediction is satisfying, as the discrepancy between target and prediction is less than 3 ms for 75% of predictions.
It is worth noting that if the specific goal is the prediction of this quantity, and if the user has a training set of APD90s for different drug configurations, a classical machine learning algorithm (that focuses on this specific quantity) might be a better option. For instance, Table 2 illustrates that a dedicated random forest performs quite well for this task. This is not a surprise as a perfect prediction of the APD90 from the voltage curves requires a perfect approximation of the lowest value, the highest value, and the decreasing rate. A small error on those three quantities could accumulate when predicting the APD90. Moreover, as shown before, the hyperSBINN model is susceptible to extreme values in the error. Having a specific sampling design for problematic drug configurations (instead of the uniform sampling chosen here) is certainly a lead for future work.
Method Min Q1 Median Mean Q3 Max
hyperSBINN 0.000 0.600 1.300 2.972 2.900 90.800
RF 0.000 0.071 0.189 0.627 0.454 25.613
@@ 7 Conclusion and perspectives
The complexity of our HyperSBINN model stems from various factors. The system involves numerous interacting equations, making the training process longer and more resource-intensive. Additionally, the model features a high number of non-fixed parameters (5 parameters that represent the drug configuration $\mathsf{X}$) with a large range of the IC50 values $([0.1-100])$. This significantly expands the search space, requiring the optimizer to explore extensively to find the optimal solution, potentially leading to slow training times. The numerous factors mentioned above contribute to a slower optimization process. The optimizer might take a long time to find a good solution or risk getting stuck in local minima. With a complex differential equation system and many parameters, there is a higher risk of the model overfitting to the training data. This means the model might perform well on the training data but fail to generalize to unseen data. The hyperparameters of the neural network become even more crucial for successful training. Careful tuning is necessary to balance exploration and exploitation during optimization.
@@
The results indicate that our model might achieve high accuracy in replicating results from traditional solvers. This close alignment underscores the model’s capability to effectively capture the dynamics of complex systems. Furthermore, the HyperSBINN model demonstrated robust performance in predicting APD90 values, showing comparable accuracy to a well-tuned Random Forest model.
@@
In our work, we leverage Data-driving learning alongside the HyperSBINN model for several key reasons. First, incorporating observations, even a small portion, significantly improves the accuracy and convergence, especially when dealing with 1024 different curve solutions during the training. This is particularly beneficial for complex systems of Ordinary Differential Equations (ODEs) where the model can encounter challenges identifying the correct solution manifold from scratch. Also, these observations act as guideposts, steering the HyperSBINN towards the physically relevant solutions within the vast solution space. Compared to literature findings where 10% to 20% of data might be used for a single curve solution, our approach shows remarkable efficiency. We achieve comparable accuracy using only 0.6% of the total observations data points for 1024 drugs configurations used for training. This is possible due to the model’s ability to learn from limited data and generalize across different configurations. This efficiency is particularly important when data acquisition is expensive.
@@
To the best of our knowledge, this is the first instance where a systems biology-informed neural network is combined with a hypernetwork for the study of cardiac action potentials. Looking forward, further refinement and validation of the HyperSBINN model on more specific and diverse datasets are necessary to fully realize its potential. Expanding the framework with real clinical and in vitro data will be crucial for translating the model advancements into pharmaceutical applications. Additionally, improving the selection of observational data ( explained in Section 5.4), ensuring that data points are more precise and representative, will enhance the model accuracy. For instance, incorporating more targeted observations at critical time points or under specific conditions could provide better guidance for the model. Furthermore, in our study, we compared our results with the solutions of the ODEs derived from a numerical solver. It would be interesting to also compare these results with real cardiac voltage curves and ionic currents derived from biological assays. Including real observation data for the gates and voltages would also be beneficial. Given that real data often contains noise, methodologies such as those proposed in the work Linka et al. (2022). Additionally, it can be interesting to adaptively manage collocation points based on model residuals, which could enhance the efficiency and accuracy of the training process as inNguyen et al. (2023).
@@ Appendix A Full cardiac action potential model
To lighten the notation, the dependence on $\mathsf{X}$ of the voltage function is omitted (i.e. $V(t)$ is written instead of $V(t,\mathsf{X})$.
A.1 Channel directly impacted by the drug characteristics
@@ A.1.1 Fast sodium current $I_{\text{Na-F}}$
$I_{\text{Na-F}}(t,\mathsf{X})$ $=\mathsf{G}_{\text{Na-F}}\times\left(1+\frac{\mathsf{D}}{\mathsf{IC50}_{\text{Na-F}}}\right)^{-1}\ \times g_{\text{Na-F}}(t)\times(V(t)-\mathsf{E}_{\text{Na-F}})\,,$
@@
$g_{\text{Na-F}}(t)$ $=f_{\text{Na-F}}(t)\times h_{\text{Na-F}}(t)$
with these two functions satisfying the following ODEs:
$\frac{{\rm{d}}f_{\text{Na-F}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(\frac{V(t)+25}{5}\right)-f_{\text{Na-F}}(t)}{0.005}\,,$
$\frac{{\rm{d}}h_{\text{Na-F}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(-\frac{V(t)+69}{3.96}\right)-h_{\text{Na-F}}(t)}{2}\,.$
@@ A.1.2 $L$-type calcium current $I_{\text{Ca}}$
$I_{\text{Ca}}(t,\mathsf{X})$ $=\mathsf{G}_{\text{Ca}}\times\left(1+\frac{\mathsf{D}}{\mathsf{IC50}_{\text{Ca}}}\right)^{-1}\ \times g_{\text{Ca}}(t)\times(V(t)-\mathsf{E}_{\text{Ca}})\,,$
@@
$g_{\text{Ca}}(t)$ $=f_{\text{Ca}}(t)\times h_{\text{Ca}}(t)$
with these two functions satisfying the following ODEs:
$\frac{{\rm{d}}f_{\text{Ca}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(\frac{V(t)+14.6}{5.5}\right)-f_{\text{Ca}}(t)}{0.7}\,,$
$\frac{{\rm{d}}h_{\text{Ca}}}{{\rm{d}}t}(t)$ $=\left(0.7\text{e}^{-0.0337\times(V(t)+14.5)^{2}}+0.04\right)\times\frac{\sigma\left(-\frac{V(t)+31}{5.54}\right)-h_{\text{Ca}}(t)}{25.1}\,.$
@@ A.1.3 Fast transient outward potassium current $I_{\text{TO-F}}$
$I_{\text{TO-F}}(t,\mathsf{X})$ $=\mathsf{G}_{\text{TO-F}}\times\left(1+\frac{\mathsf{D}}{\mathsf{IC50}_{\text{TO-F}}}\right)^{-1}\ \times g_{\text{TO-F}}(t)\times(V(t)-\mathsf{E}_{\text{K}})\,,$
@@
$g_{\text{TO-F}}(t)$ $=f_{\text{TO-F}}(t)\times h_{\text{TO-F}}(t)$
with these two functions satisfying the following ODEs:
$\frac{{\rm{d}}f_{\text{TO-F}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(\frac{V(t)+3}{15}\right)-r_{\text{TO-F}}(t)}{3.5\times\text{e}^{-\frac{V^{2}}{900}}+1.5}\,,$
$\frac{{\rm{d}}h_{\text{TO-F}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(-\frac{V(t)+33.5}{10}\right)-s_{\text{TO-F}}(t)}{20\times\left(\sigma\left(-\frac{V(t)+33.5}{10}\right)+1\right)}\,.$
@@ A.1.4 Rapidly activating delayed rectifier potassium $I_{\text{K-R}}$
$I_{\text{K-R}}(t,\mathsf{X})$ $=\mathsf{G}_{\text{K-R}}\times\left(1+\frac{\mathsf{D}}{\mathsf{IC50}_{\text{K-R}}}\right)^{-1}\ \times g_{\text{K-R}}(t)\times(V(t)-\mathsf{E}_{\text{K}})\,,$
@@
$g_{\text{K-R}}(t)$ $=f_{\text{K-R}}(t)\times\sigma\left(-\frac{V+33}{22.4}\right)$
with the function $f_{\text{K-R}}$ satisfying the following ODE:
$\frac{{\rm{d}}f_{\text{K-R}}}{{\rm{d}}t}(t)$ $=\left(\frac{0.00138\times(V(t)+7)}{1-\text{e}^{-0.123\times(V(t)+7)}}-\frac{0.00061\times(V(t)+10)}{1+\text{e}^{0.145\times(V(t)+10)}}\right)$
×$~{}~{}~{}~{}~{}~{}~{}\times\left(\sigma\left(\frac{V(t)+50}{7.5}\right)-f_{\text{K-R}}(t)\right)\,.$
A.2 Channel indirectly impacted by the drug

### v2

@@ 6.1 Performance of the hyperSBINN model
Figure 2 presents the convergence behavior of the loss components described in Section 3.1 during the training of the hyperSBINN model. The graph demonstrates the training dynamics of the hyperSBINN, with a significant initial drop in the dynamic loss ($\mathcal{L}^{\text{ODE}}$) indicating rapid learning of the underlying ODEs. The steady decrease and eventual stabilization of the initial condition loss and observation loss suggest effective learning of initial conditions and a good fit to the observed data. One can notice high variability of the ODE loss, which makes the learning highly sensitive to the learning rate schedule. The training of the hyperSBINN takes 18 minutes on a Nvidia T600 laptop GPU with the Python jinns library. Then, reconstructing the solution curves, conducted via vectorized pass-forward evaluations in the neural architecture, takes $1.8\times 10^{-2}$s. This is approximately $20$ times faster than using the ODE solver ($3.2\times 10^{-1}$s for reconstructing the curves).
@@
An example output from the hyperSBINN model on a test configuration never seen during the training (specific values are given on Appendix C) on Figure 3. One can see a close alignment for the 14 output curves, indicating that the hyperSBINN effectively manages to capture the system’s dynamics and can provide comparable performance to the traditional solver. The slight discrepancies in signal with low range of variations (such as $h_{\text{Na-L}}$ or $h_{\text{To-S}}$) might be corrected by adjustment of the different ODE weights. A comparison with the alternatives for the hyperSBINN on this same drug configuration and on another drug configuration is shown on Figures 5 and 6, Appendix C. These figures tend to show that both the hyperSBINN with the full loss and the hyperSBINN using only the observations tend to mimick well the ground truth ODE solver, while the two other alternatives provide poor results. This conjecture is confirmed by Figure 4 which provides the normalized squared differences between curves outputted by the numerical solver and curves provided by the hyperSBINN on the test set of 500 drug configurations. One can see that for most channels, the hyperSBINN trained with both the biology informed and observation losses is the best. However, the hyperSBINN learnt without the ODE loss performs systematically better on three gate functions, namely $h_{\text{K-S}},h_{\text{Ca-L}}$ and $h_{\text{To-S}}$. It is worth noting that these three gate functions are channels whose derivatives tends to be rather small (as shown on an example on Figure 3 Moreover, for three configurations, the hyperSBINN critically fails in comparison to the hyperSBINN without ODE. This figure also highlights the importance of incorporating ground truth observations in the training, as shown by the poor performances by the hyperSBINN relying only on the biology informed loss. Finally, one can see that the metaSBINN architecture leads to poorer result than the hyperSBINN architecture. This provides a new empirical evidence of the relevance of the hypernetwork.
Figure 3: Comparison of the solver and hyperSBINN for a test drug configuration never seen on the training set for each 14 solutions of the biological system. Meaning of all signals names is depicted in Appendix A. Our main signal of interest is $V(t)$ (top left). Outputs of our three alternatives to the hyperSBINN are shown in Appendix.
Figure 4: Normalized RMSE (in $\log_{10}$-scale) obtained for each of 14 channels, for 4 different models. Each boxplot is obtained thanks to 500 test curves
@@ 6.2 Learning specific biomarker of the CAP
Table 1 shows the absolute error of prediction of APD90 when this quantity is computed using the voltage curve output by the hyperSBINN. Overall, the prediction is satisfying, as the discrepancy between target and prediction is less than 3.5 ms for 75% of predictions.
It is worth noting that if the specific goal is the prediction of this quantity, and if the user has a training set of APD90s for different drug configurations, a classical machine learning algorithm (that focuses on this specific quantity) might be a better option. For instance, Table 1 illustrates that a dedicated random forest performs quite well for this task. This is not a surprise as a perfect prediction of the APD90 from the voltage curves requires a perfect approximation of the lowest value, the highest value, and the decreasing rate. A small error on those three quantities could accumulate when predicting the APD90. Moreover, the hyperSBINN model is prone to extreme values in the error. Indeed, as one can see on Figure 4 or on Figure 7 (Appendix C), two configurations lead to critically bad prediction of the voltage curve, leading to a catastrophic prediction of the APD90. Having a specific sampling design for problematic drug configurations (instead of the uniform sampling chosen here) is certainly a lead for future work.
Min Q1 Median Mean Q3 Max
HyperSBINN 0.0 0.600 1.400 4.140 3.400 173.600
HyperSBINN w/o ODE 0.1 7.400 9.050 9.899 10.925 51.600
HyperSBINN w/o obs. 299.4 310.200 311.200 313.628 313.800 394.000
MetaSBINN 0.1 9.675 15.850 16.306 22.600 65.700
Random forest 0.0 0.380 0.625 1.067 0.989 26.006
@@ 7 Conclusion and perspectives
This work shows promising results for the metamodelling of a complex biological system. The hyperSBINN achieves high accuracy in replicating results from traditional solvers for a large range of drug configurations. This close alignment underscores the model’s capability to effectively capture the dynamics of complex systems.
@@
Although based on underlying biological system, this work shows that incorporating synthetic data as observations, even a small portion, significantly improves the accuracy and convergence. This is particularly beneficial for complex systems of Ordinary Differential Equations (ODEs) where the model can encounter challenges identifying the correct solution manifold from scratch, as the poor results of the model using only the ODE loss show. These observations then act as guideposts, steering the hyperSBINN towards the physically relevant solutions within the vast solution space.
Our results also highlight the relevance of using a hypernetwork, decoupling the role of the drug configuration and time inputs in the predictive model. In addition to the better performances of the hyperSBINN framework compared to a case were time and drug inputs are equally considered, the hypernetwork modelling has a modular architecture that could ease adaptation of the model to other case study.
However, our hyperSBINN model remains complex, resulting in a training phase that can be long and sensitive to various factors. As the system involves numerous interacting equations, and multiple related loss functions, a critical tuning is the choice of loss weights. A potential improvement would be an adaptive procedure for tuning the loss weights during training, increasing the weights of loss components that have high values or show no decrease. This could help the hyperSBINN to have equal performances on all channels, as we showed here that performance over some channels (whose derivatives tend to be small) could be increased. Such procedure has been proposed by Gao et al. (2025) for partial differential equations, and would definitely help to have a more stable and easier to tune learning phase.
A second important point is the learning rate schedule which is, as often for such models, critical, as careful tuning is necessary to balance between exploration and stability during the optimization.
As discussed here, the hyperSBINN framework can be seen as ”mesh free” solver. In this work, we chose uniform sampling of collocation points both for the solving times of the ODE system, and for the drug configurations. The results highlight that, while more than often good, the prediction can show bad performances for certain drug configuration. Uniform sampling, although intuitive, might not be well suited, as it tends to show some aggregating patterns for the sample. A potential alternative is to introduce some repulsion using Sobol sampling, which would ensure a larger minimal distance between sampling points, while still offering the exploration performances of random sampling. Another attractive method would be an adaptative preferential sampling procedure, which would target configurations (or sampling times) for which the ODE loss shows high values (see Wu et al. (2023) or Bischof and Kraus (2025) and the references therein). A promising direction for future research would be to carry out a comprehensive comparison between the hyperSBINN framework and operator learning approaches (such as advocated in Huang et al. (2024), aimed at assessing the relative benefits and limitations of each method.
@@
To the best of our knowledge, this is the first instance where a system biology-informed neural network is combined with a hypernetwork for the study of cardiac action potentials. Looking forward, further refinement and validation of the hyperSBINN model on more specific and diverse datasets are necessary to fully realize its potential. Expanding the framework with real clinical and in vitro data will be crucial for translating the model advancements into pharmaceutical applications. Additionally, improving the selection of observational data could enhance the model performance for practical purposes. For instance, incorporating more targeted observations at critical time points or under specific conditions could provide better guidance for the model. Furthermore, in our study, we compared our results with the solutions of the ODEs derived from a numerical solver. It would be interesting to also compare these results with real cardiac voltage curves and ionic currents derived from biological assays. Including real observation data for the gates and voltages would also be beneficial. Given that real data often contains noise, methodologies such as those proposed in the work Linka et al. (2022) could be explored.
@@ Acknowledgement
Contribution Statement
  • I. Soukarieh: Main author, Conceptualization, Methodology, Software, Formal analysis, Writing – original draft, Investigation.
  • H. Minoux: Supervision, Project administration, Funding acquisition
  • M. Mohr, G. Hessler, J. Wenzel: Conceptualization, Methodology, Formal analysis.
  • F. Schmidt: Conceptualization, Methodology, Formal analysis, Writing-review & editing.
  • H. Gangloff: Software, Validation, Writing – review & editing
  • P.Barbillon, P. Gloaguen: Conceptualization, Methodology, Formal analysis, Validation, Writing – review & editing, Supervision
@@ Appendix A Full cardiac action potential model
The ODE system below depends on five drug characteristics denoted:
$\mathsf{X}=\left\{\mathsf{D},\mathsf{IC50}_{\text{K-R}},\mathsf{IC50}_{\text{Ca}},\mathsf{IC50}_{\text{To-F}},\mathsf{IC50}_{\text{Na-F}}\right\}\,$,
where $\mathsf{D}$ is the concentration of the drug, and the four other terms gives the inhibitory value of the drug on four ionic currents. To lighten the notation, the dependence on $\mathsf{X}$ of the voltage function is omitted, i.e. we write $V(t)$ instead of $V(t,\mathsf{X})$.
A.1 Channels directly impacted by drug configuration
@@ A.1.1 Fast sodium current $I_{\text{Na-F}}$
$I_{\text{Na-F}}(t,\mathsf{X})=\mathsf{G}_{\text{Na-F}}\times\left(1+\frac{\mathsf{D}}{\mathsf{IC50}_{\text{Na-F}}}\right)^{-1}\ \times f_{\text{Na-F}}(t)\times h_{\text{Na-F}}(t)\times(V(t)-\mathsf{E}_{\text{Na-F}})\,$,
@@
$\frac{{\rm{d}}f_{\text{Na-F}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(\frac{V(t)+25}{5}\right)-f_{\text{Na-F}}(t)}{0.005}\,$, (5)
$\frac{{\rm{d}}h_{\text{Na-F}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(-\frac{V(t)+69}{3.96}\right)-h_{\text{Na-F}}(t)}{2}\,.$ (6)
@@ A.1.2 $L$-type calcium current $I_{\text{Ca}}$
$I_{\text{Ca}}(t,\mathsf{X})=\mathsf{G}_{\text{Ca}}\times\left(1+\frac{\mathsf{D}}{\mathsf{IC50}_{\text{Ca}}}\right)^{-1}\ \times f_{\text{Ca}}(t)\times h_{\text{Ca}}(t)\times(V(t)-\mathsf{E}_{\text{Ca}})\,$,
@@
$\frac{{\rm{d}}f_{\text{Ca}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(\frac{V(t)+14.6}{5.5}\right)-f_{\text{Ca}}(t)}{0.7}\,$, (7)
$\frac{{\rm{d}}h_{\text{Ca}}}{{\rm{d}}t}(t)$ $=\left(0.7\text{e}^{-0.0337\times(V(t)+14.5)^{2}}+0.04\right)\times\frac{\sigma\left(-\frac{V(t)+31}{5.54}\right)-h_{\text{Ca}}(t)}{25.1}\,.$ (8)
@@ A.1.3 Fast transient outward potassium current $I_{\text{To-F}}$
$I_{\text{To-F}}(t,\mathsf{X})=\mathsf{G}_{\text{To-F}}\times\left(1+\frac{\mathsf{D}}{\mathsf{IC50}_{\text{To-F}}}\right)^{-1}\ \times f_{\text{To-F}}(t)\times h_{\text{To-F}}(t)\times(V(t)-\mathsf{E}_{\text{K}})\,$,
@@
$\frac{{\rm{d}}f_{\text{To-F}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(\frac{V(t)+3}{15}\right)-f_{\text{To-F}}(t)}{3.5\times\text{e}^{-\frac{V^{2}}{900}}+1.5}\,$, (9)
$\frac{{\rm{d}}h_{\text{To-F}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(-\frac{V(t)+33.5}{10}\right)-h_{\text{To-F}}(t)}{20\times\left(\sigma\left(-\frac{V(t)+33.5}{10}\right)+1\right)}\,.$ (10)
@@ A.1.4 Rapidly activating delayed rectifier potassium $I_{\text{K-R}}$
$I_{\text{K-R}}(t,\mathsf{X})=\mathsf{G}_{\text{K-R}}\times\left(1+\frac{\mathsf{D}}{\mathsf{IC50}_{\text{K-R}}}\right)^{-1}\ \times f_{\text{K-R}}(t)\times\sigma\left(-\frac{V(t)+33}{22.4}\right)\times(V(t)-\mathsf{E}_{\text{K}})\,$,
@@
$\frac{{\rm{d}}f_{\text{K-R}}}{{\rm{d}}t}(t)$ $=\left(\frac{0.00138\times(V(t)+7)}{1-\text{e}^{-0.123\times(V(t)+7)}}-\frac{0.00061\times(V(t)+10)}{1+\text{e}^{0.145\times(V(t)+10)}}\right)$ (11)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \times\left(\sigma\left(\frac{V(t)+50}{7.5}\right)-f_{\text{K-R}}(t)\right)\,.$
A.2 Channels indirectly impacted by drug configuration